# A Comprehensive Overview of Large Language Models

### v9

@@ Abstract
Large Language Models (LLMs) have recently demonstrated remarkable capabilities in natural language processing tasks and beyond. This success of LLMs has led to a large influx of research contributions in this direction. These works encompass diverse topics such as architectural innovations, better training strategies, context length improvements, fine-tuning, multi-modal LLMs, robotics, datasets, benchmarking, efficiency, and more. With the rapid development of techniques and regular breakthroughs in LLM research, it has become considerably challenging to perceive the bigger picture of the advances in this direction. Considering the rapidly emerging plethora of literature on LLMs, it is imperative that the research community is able to benefit from a concise yet comprehensive overview of the recent developments in this field. This article provides an overview of the existing literature on a broad range of LLM-related concepts. Our self-contained comprehensive overview of LLMs discusses relevant background concepts along with covering the advanced topics at the frontier of research in LLMs. This review article is intended to not only provide a systematic survey but also a quick comprehensive reference for the researchers and practitioners to draw insights from extensive informative summaries of the existing works to advance the LLM research.
@@
Figure 1: The trend of papers released over years containing keywords "Large Language Model", "Large Language Model + Fine-Tuning", and "Large Language Model + Alignment".
Figure 2: Chronological display of LLM releases: blue cards represent ‘pre-trained’ models, while orange cards correspond to ‘instruction-tuned’ models. Models on the upper half signify open-source availability, whereas those on the bottom half are closed-source. The chart illustrates the increasing trend towards instruction-tuned models and open-source models, highlighting the evolving landscape and trends in natural language processing research.
@@ 1 Introduction
Language plays a fundamental role in facilitating communication and self-expression for humans, and their interaction with machines. The need for generalized models stems from the growing demand for machines to handle complex language tasks, including translation, summarization, information retrieval, conversational interactions, etc. Recently, significant breakthroughs have been witnessed in language models, primarily attributed to transformers [1], increased computational capabilities, and the availability of large-scale training data. These developments have brought about a revolutionary transformation by enabling the creation of LLMs that can approximate human-level performance on various tasks [2, 3]. Large Language Models (LLMs) have emerged as cutting-edge artificial intelligence systems that can process and generate text with coherent communication [4], and generalize to multiple tasks [5, 6].
@@
In addition to better generalization and domain adaptation, LLMs appear to have emergent abilities, such as reasoning, planning, decision-making, in-context learning, answering in zero-shot settings, etc. These abilities are known to be acquired by them due to their gigantic scale even when the pre-trained LLMs are not trained specifically to possess these attributes [22, 23, 24]. Such abilities have led LLMs to be widely adopted in diverse settings including, multi-modal, robotics, tool manipulation, question answering, autonomous agents, etc. Various improvements have also been suggested in these areas either by task-specific training [25, 26, 27, 28, 29, 30, 31] or better prompting [32].
@@
The LLMs abilities to solve diverse tasks with human-level performance come at a cost of slow training and inference, extensive hardware requirements, and higher running costs. Such requirements have limited their adoption and opened up opportunities to devise better architectures [15, 33, 34, 35] and training strategies [36, 37, 21, 38, 39, 40, 41]. Parameter efficient tuning [38, 41, 40], pruning [42, 43], quantization [44, 45], knowledge distillation, and context length interpolation [46, 47, 48, 49] among others are some of the methods widely studied for efficient LLM utilization.
@@
    We present a survey on the developments in LLM research providing a concise comprehensive overview of the direction.
@@ 3.1.1 General Purpose
[…]
@@
[…]
@@
 UL2 [125]: An encoder-decoder architecture trained using a mixture of denoisers (MoD) objective. Denoisers include 1) R-Denoiser: a regular span masking, 2) S-Denoiser: which corrupts consecutive tokens of a large sequence and 3) X-Denoiser: which corrupts a large number of tokens randomly. During pre-training, UL2 includes a denoiser token from R,S,X{R,S,X} to represent a denoising setup. It helps improve fine-tuning performance for downstream tasks that bind the task to one of the upstream training modes. This MoD style of training outperforms the T5 model on many benchmarks.
@@
 PanGu-$\Sigma$ [92]: An autoregressive model with parameters copied from PanGu-$\alpha$ and extended to a trillion scale with Random Routed Experts (RRE), the architectural diagram is shown in Figure 10. RRE is similar to the MoE architecture, with distinctions at the second level, where tokens are randomly routed to experts in a domain instead of using a learnable gating method. The model has bottom layers densely activated and shared across all domains, whereas top layers are sparsely activated according to the domain. This training style allows extracting task-specific models and reduces catastrophic forgetting effects in the case of continual learning.
@@ 3.1.2 Coding
 CodeGen [130]: CodeGen has similar architecture to PaLM [15], i.e., parallel attention, MLP layers, and RoPE embeddings. The model is trained on both natural language and programming language data sequentially (trained on the first dataset, then the second and so on) on the following datasets 1) PILE, 2) BIGQUERY and 3) BIGPYTHON. CodeGen proposed a multi-step approach to synthesizing code. The purpose is to simplify the generation of long sequences where the previous prompt and generated code are given as input with the next prompt to generate the next code sequence. CodeGen opensource a Multi-Turn Programming Benchmark (MTPB) to evaluate multi-step program synthesis.
 Codex [131]: This LLM is trained on a subset of public Python Github repositories to generate code from docstrings. Computer programming is an iterative process where the programs are often debugged and updated before fulfilling the requirements. Similarly to this, Codex generates 100 versions of a program by repetitive sampling for a given description, which produces a working solution for 77.5% of the problems passing unit tests. Its powerful version powers Github Copilot [^2] https://github.com/features/copilot.
@@ 3.1.3 Scientific Knowledge
 Galactica [138]: A large curated corpus of human scientific knowledge with 48 million papers, textbooks, lecture notes, millions of compounds and proteins, scientific websites, encyclopedias, and more are trained using the metaseq library3, which is built on PyTorch and fairscale [139]. The model wraps reasoning datasets with the <work><work> token to provide step-by-step reasoning context to the model, which has been shown to improve the performance on reasoning tasks.
@@ 3.1.5 Finance
 BloombergGPT [141]: A non-causal decoder model trained using both financial ("FINPILE" from the Bloomberg archive) and general-purpose datasets. The model’s architecture is similar to the BLOOM [13] and OPT [14]. It allocates 50B parameters to different blocks of the model using the approach [113]. For effective training, BloombergGPT packs documents together with <|endoftext|><|endoftext|> to use the maximum sequence length, uses warmup batch size starting from 1024 to 2048, and manually reduces the learning rate multiple times during the training.
@@ 1. LLMs have good few shot capabilities
1. Parallel attention + $F$ layers speed-up training 15% with the same performance as with cascaded layers 2. Initializing feed-forward output layers before residuals with scheme in [143] avoids activations from growing with increasing depth and width 3. Training on Pile outperforms GPT-3 on five-shot
@@ 3.4.2 Tool Augmented LLMs
Multimodal Tool Augmentation: The compositional reasoning capacity of LLMs allows them to manipulate tools in multimodal settings [205, 206, 214]. Following the pipeline shown in Figure 13, the LLM outlines a plan, generally executing in a sequence: Plan $\to\,$ Tool selection $\to\,$ Execute $\to\,$ Inspect $\to\,$ Generate, to respond to the user query. Here, the database of tools is rich in modalities, including text, images, etc. Many of the multimodal tool augmentation systems employ multimodal LLMs [31, 215, 214, 206], while others utilize single modality LLMs and generate a plan on using different modality tools to solve multimodal queries [216].
@@ 3.6.2 Quantization
LLMs require extensive computing and memory for inference. Deploying a 175B parameter GPT-3 model needs at least 5x80GB A100 GPUs and 350GB of memory to store in FP16 format [44]. Such demanding requirements for deploying LLMs make it harder for smaller organizations to utilize them. Model compression is an effective solution but comes at the cost of degraded performance, especially at large scales greater than 6B. These models exhibit very large magnitude outliers that do not exist in smaller models [245], making it challenging and requiring specialized methods for quantizing LLMs [44, 246].
@@ 3.8.5 Encoder vs Decoder vs Encoder-Decoder
Table 3: Summary of pre-trained LLMs (>10B). Only the LLMs discussed individually in the previous sections are summarized. “Data/Tokens” is the model’s pre-training data, which is either the number of tokens or data size. “Data Cleaning” indicates whether data cleaning is performed or not. This includes heuristics (Heur), deduplication (Dedup), quality filtering (QF), and privacy filtering (PF), “Cost” is the calculated training cost obtained by multiplying the GPUs/TPUs hourly rate with the number of GPUs and the training time. The actual cost may vary due to many reasons such as using in-house GPUs or getting a discounted rate, re-training, number of employees working on the problem, etc. “Training Parallelism” indicates distributed training using data parallelism (D), tensor parallelism (T), pipeline parallelism (P), model parallelism (M), optimizer parallelism (OP), and rematerialization (R), where for “Library” column, “DS” is a short form for Deep Speed. In column “Commercial Use”, we assumed a model is for non-commercial purposes if its license is unavailable.
@@
Xuan Yuan 2.0 [142] arXiv23 RAIL-1.0 Du Xiaoman Finance 176B $\checkmark$ - 366B Filtered 80GB A100 - - P DS
@@
Table 4: Summary of instruction tuned LLMs (>10B). All abbreviations are the same as Table 3. Entries in “Data/Tokens” starting with “S-” represents the number of training samples.
@@ 4 Model Configurations
LLaMA (70B) Causal-Dec Next Token Grouped-query 32k BPE Pre-RMS RoPE SwiGLUE - - - -
@@
Yuan 1.0 (245B) <<10M 2048 1.6e-4 $\checkmark$ cosine decay to 10% $\checkmark$ - - - $\checkmark$ - -

### v10

@@ Abstract
Large Language Models (LLMs) have recently demonstrated remarkable capabilities in natural language processing tasks and beyond. This success of LLMs has led to a large influx of research contributions in this direction. These works encompass diverse topics such as architectural innovations, better training strategies, context length improvements, fine-tuning, multi-modal LLMs, robotics, datasets, benchmarking, efficiency, and more. With the rapid development of techniques and regular breakthroughs in LLM research, it has become considerably challenging to perceive the bigger picture of the advances in this direction. Considering the rapidly emerging plethora of literature on LLMs, it is imperative that the research community is able to benefit from a concise yet comprehensive overview of the recent developments in this field. This article provides an overview of the literature on a broad range of LLM-related concepts. Our self-contained comprehensive overview of LLMs discusses relevant background concepts along with covering the advanced topics at the frontier of research in LLMs. This review article is intended to provide not only a systematic survey but also a quick, comprehensive reference for the researchers and practitioners to draw insights from extensive, informative summaries of the existing works to advance the LLM research.
@@
Figure 1: The trend of papers released over the years containing keywords "Large Language Model", "Large Language Model + Fine-Tuning", and "Large Language Model + Alignment".
Figure 2: Chronological display of LLM releases: blue cards represent ‘pre-trained’ models, while orange cards correspond to ‘instruction-tuned’ models. Models on the upper half signify open-source availability, whereas those on the bottom are closed-source. The chart illustrates the increasing trend towards instruction-tuned and open-source models, highlighting the evolving landscape and trends in natural language processing research.
@@ 1 Introduction
Language plays a fundamental role in facilitating communication and self-expression for humans and their interaction with machines. The need for generalized models stems from the growing demand for machines to handle complex language tasks, including translation, summarization, information retrieval, conversational interactions, etc. Recently, significant breakthroughs have been witnessed in language models, primarily attributed to transformers [1], increased computational capabilities, and the availability of large-scale training data. These developments have brought about a revolutionary transformation by enabling the creation of LLMs that can approximate human-level performance on various tasks [2, 3]. Large Language Models (LLMs) have emerged as cutting-edge artificial intelligence systems that can process and generate text with coherent communication [4] and generalize to multiple tasks [5, 6].
@@
In addition to better generalization and domain adaptation, LLMs appear to have emergent abilities, such as reasoning, planning, decision-making, in-context learning, answering in zero-shot settings, etc. These abilities are known to be acquired by them due to their gigantic scale even when the pre-trained LLMs are not trained specifically to possess these attributes [22, 23, 24]. Such abilities have led LLMs to be widely adopted in diverse settings, including multi-modal, robotics, tool manipulation, question answering, autonomous agents, etc. Various improvements have also been suggested in these areas either by task-specific training [25, 26, 27, 28, 29, 30, 31] or better prompting [32].
@@
The LLMs abilities to solve diverse tasks with human-level performance come at the cost of slow training and inference, extensive hardware requirements, and higher running costs. Such requirements have limited their adoption and opened up opportunities to devise better architectures [15, 33, 34, 35] and training strategies [36, 37, 21, 38, 39, 40, 41]. Parameter efficient tuning [38, 41, 40], pruning [42, 43], quantization [44, 45], knowledge distillation, and context length interpolation [46, 47, 48, 49] among others are some of the methods widely studied for efficient LLM utilization.
@@
    We present a survey on the developments in LLM research, providing a concise, comprehensive overview of the direction.
@@ 3.1.1 General Purpose
 GLaM [91]: Generalist Language Model (GLaM) represents a family of language models using a sparsely activated decoder-only mixture-of-experts (MoE) structure [121, 90]. To gain more model capacity while reducing computation, the experts are sparsely activated where only the best two experts are used to process each input token. The largest GLaM model, GLaM (64B/64E), is about 7$\times$ larger than GPT-3 [6], while only part of the parameters are activated per input token. The largest GLaM (64B/64E) model achieves better overall results as compared to GPT-3 while consuming only one-third of GPT-3’s training energy.
@@
 MT-NLG [117]: A 530B causal decoder based on the GPT-2 architecture that has roughly 3$\times$ GPT-3 model parameters. MT-NLG is trained on filtered high-quality data collected from various public datasets and blends various types of datasets in a single batch, which beats GPT-3 on several evaluations.
@@
 UL2 [125]: An encoder-decoder architecture trained using a mixture of denoisers (MoD) objective. Denoisers include 1) R-Denoiser: a regular span masking, 2) S-Denoiser: which corrupts consecutive tokens of a large sequence and 3) X-Denoiser: which corrupts a large number of tokens randomly. During pre-training, UL2 includes a denoiser token from ${R,S,X}$ to represent a denoising setup. It helps improve fine-tuning performance for downstream tasks that bind the task to one of the upstream training modes. This MoD style of training outperforms the T5 model on many benchmarks.
@@
 LLaMA-3/3.1 [130]: A collection of models trained on a seven times larger dataset as compared to LLaMA-2 with double the context length, outperforming its previous variants and other models.
 PanGu-$\Sigma$ [92]: An autoregressive model with parameters copied from PanGu-$\alpha$ and extended to a trillion scale with Random Routed Experts (RRE), the architectural diagram is shown in Figure 10. RRE is similar to the MoE architecture, with distinctions at the second level, where tokens are randomly routed to experts in a domain instead of using a learnable gating method. The model has bottom layers densely activated and shared across all domains, whereas top layers are sparsely activated according to the domain. This training style allows for extracting task-specific models and reduces catastrophic forgetting effects in the case of continual learning.
 Mixtral8x22b [131]: A mixture-of-experts (MoE) model with eight distinct experts routes each token to two experts at each layer and combines the outputs additively.
 Snowflake Arctic [132]: Arctic LLM is a hybrid of dense and mixture-of-experts (MoE) architecture. The MoE (128$\times$3.66B MLP experts) is parallel to the dense transformer (10B) with only two experts activated. The model has many experts, compared to other MoE LLMs [131, 133], to increase the model capacity and provide an opportunity to choose among many experts for a diverse configuration. The model has 480B parameters, and only 17B are active during a forward pass, reducing the computation significantly.
 Grok [133, 134]: Grok is a family of LLMs including Grok-1 and Grok-1.5, released by XAI.
 Grok-1 [133]: Grok-1 is a 314B parameters language MoE model (eight experts), where two experts are activated per token.
 Grok-1.5 [134]: Grok-1.5 is a multi-modal LLM with a larger context length and improved performance.
 Gemini [135, 136]: Gemini replaces Bard (based on PaLM) with multi-modal capabilities and significant language modeling performance improvements.
 Gemini-1 [135]: The first-ever auto-regressive model to achieve human-level capabilities on the MMLU benchmark.
 Gemini-1.5 [136]: A multi-modal LLM with MoE architecture builds on the findings of Gemini-1. The model has a 2M context window and can reason over information up to 10M tokens. Such large context windows were never achieved previously and shown to have a huge impact on performance gain.
 Nemotron-4 340B [137]: A decoder-only model that has been aligned on 98% synthetic data and only 2% manually annotated data. Utilizing synthetic data at a large proportion improves the model performance significantly. The paper suggested introducing alignment data with a smaller subset of previously seen data during the late stage of the model pre-training, enabling the smooth transition from the pre-trained stage to the final training stage. To train better instruction-following models, weaker models are trained into stronger models iteratively. The synthetic data generated by the weaker instruction-tuned model is used to train a base model which is later supervised fine-tuned outperforming the weaker model.
 DeepSeek [138]: DeepSeek studies the LLMs scaling laws in detail to determine the optimal non-embedding model size and training data. The experiments were performed for 8 budgets ranging from 1e17 to 3e20 training FLOPs. Each compute budget was tested against ten different models/data scales. The batch size and learning rates were also fitted for the given compute budget finding that the batch size should increase with the increased compute budget while decreasing the learning rate. Following are the equations for the optimal batch-size ($B$), learning rate ($\eta$), model size ($M$), and data ($D$):
$\begin{gathered}B_{opt}=0.2920.C^{0.3271}\\ \eta_{opt}=0.3118.C^{-0.1250}\\ M_{opt}=M_{base}.C^{a}\\ D_{opt}=D_{base}.C^{b}\\ M_{base}=0.1715,D_{base}=5.8316,a=0.5243,b=0.4757\\ \end{gathered}$ (5)
 DeepSeek-v2 [139]: An MoE model that introduces multi-head latent attention (MLA) to reduce inference costs, by compressing Key-Value (KV) cache into a latent vector. MLA achieves better performance than multi-head attention (MHA), and other efficient attention mechanisms such as grouped query attention (GQA), multi-query attention (MQA), etc. Because of MLA, DeepSeek-v2 achieves 5.76 times faster inference throughput as compared to DeepSeek [138].
@@ 3.1.2 Coding
 CodeGen [140]: CodeGen has a similar architecture to PaLM [15], i.e., parallel attention, MLP layers, and RoPE embeddings. The model is trained on both natural language and programming language data sequentially (trained on the first dataset, then the second, and so on) on the following datasets 1) PILE, 2) BIGQUERY, and 3) BIGPYTHON. CodeGen proposed a multi-step approach to synthesizing code. The purpose is to simplify the generation of long sequences where the previous prompt and generated code are given as input with the next prompt to generate the next code sequence. CodeGen opensource a Multi-Turn Programming Benchmark (MTPB) to evaluate multi-step program synthesis.
 Codex [141]: This LLM is trained on a subset of public Python Github repositories to generate code from docstrings. Computer programming is an iterative process where the programs are often debugged and updated before fulfilling the requirements. Similarly, Codex generates 100 versions of a program by repetitive sampling for a given description, which produces a working solution for 77.5% of the problems passing unit tests. Its powerful version powers Github Copilot [^2] https://github.com/features/copilot.
@@ 3.1.3 Scientific Knowledge
 Galactica [148]: A large curated corpus of human scientific knowledge with 48 million papers, textbooks, lecture notes, millions of compounds and proteins, scientific websites, encyclopedias, and more are trained using the metaseq library3, which is built on PyTorch and fairscale [149]. The model wraps reasoning datasets with the $<work>$ token to provide step-by-step reasoning context to the model, which has been shown to improve the performance on reasoning tasks.
@@ 3.1.5 Finance
 BloombergGPT [151]: A non-causal decoder model trained using both financial (“FINPILE” from the Bloomberg archive) and general-purpose datasets. The model’s architecture is similar to the BLOOM [13] and OPT [14]. It allocates 50B parameters to different blocks of the model using the approach [113]. For effective training, BloombergGPT packs documents together with $<|endoftext|>$ to use the maximum sequence length, uses warmup batch size starting from 1024 to 2048, and manually reduces the learning rate multiple times during the training.
@@ 1. LLMs have good few shot capabilities
1. Parallel attention + FF layers speed-up training 15% with the same performance as with cascaded layers 2. Initializing feed-forward output layers before residuals with scheme in [153] avoids activations from growing with increasing depth and width 3. Training on Pile outperforms GPT-3 on five-shot
@@ 1. HHH prompt by Anthropic allows the model to follow instructions without fine-tuning
LLaMA-3/3.1
1. Increasing batch size gradually stabilizes the training without loss spikes 2. High-quality data at the final stages of training improves the model performance 3. Increasing model context length windows step-wise allows it to better adapt to various sequence lengths
Nemotron-40B
1. Model aligned iteratively on synthetic data with data generated from the previously aligned model achieves competitive performance
DeepSeek
1. Batch size should increase with the increase in compute budget while decreasing the learning rate
DeepSeek-v2
1. Mult-head latent attention (MLA) performs better than multi-head attention (MHA) while requiring a significantly smaller KV cache, therefore achieving faster data generation
@@ 3.4.2 Tool Augmented LLMs
[…]
@@ 3.6.2 Quantization
LLMs require extensive computing and memory for inference. Deploying a 175B parameter GPT-3 model needs at least five 80GB A100 GPUs and 350GB of memory to store in FP16 format [44]. Such demanding requirements for deploying LLMs make it harder for smaller organizations to utilize them. Model compression is an effective solution but comes at the cost of degraded performance, especially at large scales greater than 6B. These models exhibit very large magnitude outliers that do not exist in smaller models [255], making it challenging and requiring specialized methods for quantizing LLMs [44, 256].
@@ 3.8.5 Encoder vs Decoder vs Encoder-Decoder
Table 3: Summary of pre-trained LLMs (>10B). Only the LLMs discussed individually in the previous sections are summarized. “Data/Tokens” is the model’s pre-training data, which is either the number of tokens or data size. “Data Cleaning” indicates whether data cleaning is performed or not. This includes heuristics (Heur), deduplication (Dedup), quality filtering (QF), and privacy filtering (PF), “Cost” is the calculated training cost obtained by multiplying the GPUs/TPUs hourly rate with the number of GPUs and the training time. The actual cost may vary due to many reasons such as using in-house GPUs or getting a discounted rate, re-training, number of employees working on the problem, etc. “Training Parallelism” indicates distributed training using data parallelism (D), tensor parallelism (T), pipeline parallelism (P), context parallelism (C), model parallelism (M), optimizer parallelism (OP), and rematerialization (R), where for “Library” column, “DS” is a short form for Deep Speed. In column “Commercial Use”, we assumed a model is for non-commercial purposes if its license is unavailable.
@@
Xuan Yuan 2.0 [152] arXiv23 RAIL-1.0 Du Xiaoman Finance 176B $\checkmark$ - 366B Filtered - 80GB A100 - - P DS
@@
LLaMA-3.1 [130] arXiv'24 LLaMA-3.0 Meta General 405B $\checkmark$ 1.2M 15T Dedup+QF 16k 80G H100 30.84Mh - D+T+P+C PyTorch
Mixtral 8x22B [131] web'24 Apache-2.0 Mistral AI General 141B $\checkmark$ - - - - - - - - -
Snowflake Arctic [132] web'24 Apache-2.0 Snowflake General 480B $\checkmark$ - 3.5T - - - - T+P DS
Nemotron-4 340B [137] web'24 Nvidia Nvidia General 340B $\checkmark$ - 9T - 6144 80G H100 - - D+T+P -
DeepSeek [138] arXiv'24 MIT DeepSeek General 67B $\checkmark$ - 2T Dedup+QF - - 300.6Kh - D+T+P DS
DeepSeek-v2 [139] arXiv'24 MIT DeepSeek General 67B $\checkmark$ - 8.1T QF - H800 172.8Kh - D+P HAI-LLM
Table 4: Summary of instruction tuned LLMs (>10B). All abbreviations are the same as Table 3. Entries in “Data/Tokens” starting with “S-” represent the number of training samples.
@@ 4 Model Configurations
LLaMA-2 (70B) Causal-Dec Next Token Grouped-query 32k BPE Pre-RMS RoPE SwiGLUE - - - -
@@
LLaMA-3.1 (405B) Causal-Dec Next Token Grouped-query 128k BPE Pre-RMS RoPE SwiGLU - 126 128 16384
Nemotron-4 (340B) Causal-Dec Next Token Standard 256k SentencePiece - RoPE ReLU $\times$ 96 96 18432
DeepSeek (67B) Causal-Dec Next Token Grouped-query 100k BBPE Pre-RMS RoPE SwiGLU - 95 64 8192
DeepSeek-v2 (67B) MoE-Dec Next Token Multi-Head Latent 100k BBPE Pre-RMS RoPE SwiGLU - 60 128 5120
@@
Yuan 1.0 (245B) $<$10M 2048 1.6e-4 $\checkmark$ cosine decay to 10% $\checkmark$ - - - $\checkmark$ - -
@@
LLaMA-3.1 (405B) 16M 8192 8e-5 $\checkmark$ linear+cosine $\checkmark$ $\checkmark$ - - -
Nemotron-4 (340B) 2304 4096 - - linear - - - $\checkmark$ - - $\times$
DeepSeek (67B) 4608 4096 3.2e-4 $\checkmark$ cosine $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ -
DeepSeek-v2 (67B) 9216 4k 2.4e-4 $\checkmark$ step-decay $\checkmark$ - - - $\checkmark$ $\checkmark$ -